On June 30, 2021, Georgia-Atlantic, Inc. leased warehouse equipment from Builders, Inc. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $583,573 over a 4-year lease term (also the asset’s useful life), payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 11.0%, the same rate Builders used to calculate lease payment amounts. Builders manufactured the equipment at a cost of $3.4 million.

Required:
1. Determine the price at which Builders is “selling” the equipment (present value of the lease payments) at June 30, 2021.
2. What amount related to the lease would Builders report in its balance sheet at December 31, 2021 (ignore taxes)?
3. What line item amounts related to the lease would Builders report in its income statement for the year ended December 31, 2021
Answer #1

olution 1:
Semiannual lease payment to be received = $583,573
Total semiannual payments to be received = 4*2 = 8
Incremental borrowing rate = 11%, 5.5% semiannual
Price at which Builders is “selling” the equipment = Semi Annual lease payments * Cumulative PV Factor of annuity due for 8 periods at 5.5%
= $583573 * 6.68297 = $3,900,000

Solution 2:
Semiannual payment received on 30.06.2021 = $583,573
Pretax amount of liability on 30.06.2021 = ($3,900,000 - $583573) = $3,316,427
Interest revenue for 31.12.2021 = $3,316,427 * 5.5% = $182,403
Semiannual lease payment received on 31.12.2021 = $583,573
Pre tax amount for receivables on December 31, 2021 = $3,316,427 + $182403 - $583573 = $2,915,257

Solution 3:
Pre tax amount of interest Revenue Builder Inc.  reports in its income statement = $3316427 * 5.5% = $182,403

Pre tax amount of revenue Builder Inc.  reports in its income statement = $3,900,000

Pre tax amount of cost of goods sold to be reported by Builder Inc. in its income statement = $3,400,000

Pretax amounts related to the lease would Builders report in its income statement for the year ended December 31, 2021 = $182,403 + $3,900,000 - $3,400,000 = $682,403